# Asymptotically tight Lagrangian dual of smooth nonconvex problems via improved error bound of Shapley-Folkman Lemma

### v1

@@ 5 Asymptotically tight Lagrangian dual of smooth nonconvex problems
For any $L$-smooth function convex $f(\cdot)$ and any $\mathbf{x}\in\mathbb{R}^{n}$, there exists a $l_{2}$-ball $B$ with radius $\frac{1}{L}$ such that $B\subseteq\operatorname{\operatorname{epi}}f$ and $(\mathbf{x},f(\mathbf{x}))\subseteq B$.
Proof 5.29.
For any $\mathbf{x}\in\mathbb{R}^{n}$, let $\mathbf{g}=\nabla f(\mathbf{x})$. We construct the center of the ball $B$ with radius $r:=\frac{1}{L}$ by
$(\mathbf{x}_{c},t_{c})=(\mathbf{x},f(\mathbf{x})+\frac{1}{L}\dfrac{1}{\sqrt{\left\lVert\mathbf{g}\right\rVert_{2}^{2}+1}}\left(-\mathbf{g},1\right).$
Clearly, $(\mathbf{x},f(\mathbf{x}))\in B$ and it remains to prove that $B\subseteq\operatorname{\operatorname{epi}}f$. For every point $(\mathbf{y},t)\subseteq B$, the lowest possible $t$ is
$t_{\inf}(\mathbf{y}):=t_{c}-\sqrt{r^{2}-\left\lVert\mathbf{y}-\mathbf{x}_{c})\right\rVert_{2}^{2}}$
Therefore, it suffices to prove $f(\mathbf{y})\leq t_{\inf}(\mathbf{y})$ for all $\mathbf{y}$ such that $(\mathbf{y}-\mathbf{x}_{c})^{2}\leq r^{2}$. Note that $f(\mathbf{y})\leq t_{\inf}(\mathbf{y})$ can be further simplified to prove that
@@
For every extreme point $\mathbf{v}$ (0-dimensional face) of ${\mathbf{conv}}(\mathcal{P}^{(i)})$, there exists a ball $l_{2}$-ball $B$ with radius $\frac{\mathcal{L}_{s}(B^{(i)})}{L}$ such that $\mathbf{v}\in B\subseteq\mathcal{P}^{(i)}$.
Proof 5.30.
Fixed any support set $S\subseteq[n]$, we define
@@
$\mathcal{P}^{(i)}=\bigcup_{S\in\binom{[n]}{s_{i}}}\mathcal{P}^{(i)}_{S}$.
Let $\mathbf{v}:=(t_{\mathbf{v}},\mathbf{d}_{\mathbf{v}})$ be any extreme point of ${\mathbf{conv}}(\mathcal{P}^{(i)})$. In this case, there must exist some $S\subseteq\binom{[n]}{s_{i}}$ such that $\mathbf{v}\in P^{(i)}_{S}$. By construction of $P^{(i)}_{S}$, there exists some $\mathbf{x}_{\mathbf{v}}\in\mathbb{R}^{n}$ such that
@@
Since $\mathbf{v}$ is a extreme point, we can further assume that $t_{\mathbf{v}}=f^{(i)}(\mathbf{x})$. Consider the following function $f^{(i)}_{S}(\mathbf{y}):\mathbb{R}^{S}\to\mathbb{R}$ which is constructed from $f^{(i)}(\mathbf{x})$ by fixing the variables outside of $S$ to be zero. Clearly, $f^{(i)}_{S}$ is a $L$-smooth convex function. Applying Claim 3, there is a ball $B_{S}$ with radius $\frac{1}{L}$ with dimension $|S|+1$ that $((\mathbf{x}_{\mathbf{v}})_{S},f(\mathbf{x}_{\mathbf{v}}))\subseteq B_{S}\subseteq\operatorname{\operatorname{epi}}f_{S}^{(i)}$. Now consider the linear map:
$(\mathbf{x}_{S},t)\to\begin{bmatrix}&\textbf{1}^{\top}\\ &B_{S}^{(i)}\end{bmatrix}(\mathbf{x}_{S},t)$
Let $Q:=\begin{bmatrix}&\textbf{1}^{\top}\\ &B_{S}^{(i)}\end{bmatrix}$. It is well known that a linear map transforms a ball into an ellipse. In particular, $B_{S}$ is mapped to an ellipse $E$ (up to translation) with form: $\left\{\mathbf{y}:\mathbf{y}^{\top}(QQ^{\top})^{-1}\mathbf{y}\leq(\dfrac{1}{L})^{2}\right\}$. This ellipse includes a ball with radius $\frac{\sigma_{\inf}(QQ^{\top})}{L}$ and by the definition of projection factor, this ellipse includes a ball with radius $\frac{\mathcal{L}_{s}(B^{(i)})}{L}$. Since $E\subseteq\mathcal{P}^{(i)}_{S}\subseteq\mathcal{P}^{(i)}$, this implies that there exists a ball with the desired statement.
@@
The smoothness assumption in Theorem 5.31 can be relaxed. Instead of requiring that $f$ be globally smooth—i.e., globally upper bounded by a convex quadratic function—it suffices to assume that $f$ is locally upper bounded by a convex quadratic function at each point $\mathbf{x}$. In particular, as long as this local smoothness condition guarantees the existence of an $l_{2}$ -ball with nontrivial radius, the conclusion of Theorem 5.31 remains valid.
@@ Acknowledgments
We would also like to thank Diego Cifuentes and Greg Blekherman for the various wonderful discussions.

### v2

@@ 5 Asymptotically tight Lagrangian dual of smooth nonconvex problems
For any $L$-smooth function $f(\cdot)$ and any $\mathbf{x}\in\mathbb{R}^{n}$, there exists a $l_{2}$-ball $\mathcal{B}$ with radius $\frac{1}{L}$ such that $\mathcal{B}\subseteq\operatorname{\operatorname{epi}}f$ and $(\mathbf{x},f(\mathbf{x}))\subseteq\mathcal{B}$.
Proof 5.35.
For any $\mathbf{x}\in\mathbb{R}^{n}$, let $\mathbf{g}=\nabla f(\mathbf{x})$. We construct the center of the ball $\mathcal{B}$ with radius $r:=\frac{1}{L}$ by
$(\mathbf{x}_{c},t_{c})=(\mathbf{x},f(\mathbf{x}))+\frac{1}{L}\dfrac{1}{\sqrt{\left\lVert\mathbf{g}\right\rVert_{2}^{2}+1}}\left(-\mathbf{g},1\right).$
Clearly, $(\mathbf{x},f(\mathbf{x}))\in\mathcal{B}$ and it remains to prove that $\mathcal{B}\subseteq\operatorname{\operatorname{epi}}f$. For every point $(\mathbf{y},t)\subseteq\mathcal{B}$, the lowest possible value of $t$ is
$t_{\inf}(\mathbf{y}):=t_{c}-\sqrt{r^{2}-\left\lVert\mathbf{y}-\mathbf{x}_{c}\right\rVert_{2}^{2}}.$
Therefore, it suffices to prove $f(\mathbf{y})\leq t_{\inf}(\mathbf{y})$ for all $\mathbf{y}$ such that $\left\lVert\mathbf{y}-\mathbf{x}_{c}\right\rVert_{2}^{2}\leq r^{2}$. Note that $f(\mathbf{y})\leq t_{\inf}(\mathbf{y})$ can be further simplified to prove that
@@
Lemma 5.36.
[27] Let $E:=\{\mathbf{x}:\left\lVert Ax-\mathbf{k}\right\rVert_{2}\leq 1\}$ be an ellipsoid, where $A$ is an invertible matrix. Then, for every $\mathbf{p}\in E$, there exists a $l_{2}$-norm ball containing $\mathbf{p}$ that is subset of $E$ and has a radius of $\frac{\sqrt{\sigma_{\min}(A)}}{\sigma_{\max}(A)}$.
@@
For every extreme point $\mathbf{v}$ (0-dimensional face) of ${\mathbf{conv}}\left(\mathcal{P}^{(i)}\right)$, there exists a $l_{2}$-ball $B$ with radius $\frac{\mathcal{L}_{s}(B^{(i)})}{L}$ such that $\mathbf{v}\in B\subseteq\mathcal{P}^{(i)}$.
Proof 5.37.
For a fixed support set $S\subseteq[n]$, we define
@@
$\mathcal{P}^{(i)}=\bigcup_{S\subseteq[n],|S|=s_{i}}\mathcal{P}^{(i)}_{S}$.
Let $\mathbf{v}:=(t_{\mathbf{v}},\mathbf{d}_{\mathbf{v}})$ be any extreme point of ${\mathbf{conv}}(\mathcal{P}^{(i)})$. In this case, there must exist some $S\subseteq[n]$, with $|S|=s_{i}$, such that $\mathbf{v}\in P^{(i)}_{S}$. By construction of $P^{(i)}_{S}$, there exists some $\mathbf{x}_{\mathbf{v}}\in\mathbb{R}^{n}$ such that
@@
Since $\mathbf{v}$ is a extreme point, we can further assume that $t_{\mathbf{v}}=f^{(i)}(\mathbf{x})$. Consider the following function $f^{(i)}_{S}(\mathbf{y}):\mathbb{R}^{S}\to\mathbb{R}$ which is constructed from $f^{(i)}(\mathbf{x})$ by fixing the variables outside of $S$ to be zero. Clearly, $f^{(i)}_{S}$ is a $L$-smooth convex function. Applying Claim 5, there is a ball $\mathcal{B}_{S}$ with radius $\frac{1}{L}$ with dimension $|S|+1$ that $((\mathbf{x}_{\mathbf{v}})_{S},f(\mathbf{x}_{\mathbf{v}}))\subseteq\mathcal{B}_{S}\subseteq\operatorname{\operatorname{epi}}f_{S}^{(i)}$. Now consider the linear map:
$(\mathbf{x}_{S},t)\to\begin{bmatrix}1&\textbf{0}\\ \textbf{0}&B_{S}^{(i)}\end{bmatrix}(\mathbf{x}_{S},t)$
Let $Q:=\begin{bmatrix}1&\textbf{0}\\ \textbf{0}&B_{S}^{(i)}\end{bmatrix}$. If $QQ^{\top}$ is not invertible, we have $\mathcal{L}_{s}(B^{(i)})=0$ and the lemma directly holds. Therefore, we assume that $QQ^{\top}$ is invertible. It is well known that a linear map transforms a ball into an ellipse. In particular, $\mathcal{B}_{S}$ is mapped to an ellipse $E$ (up to translation) with form: $\left\{\mathbf{y}:\mathbf{y}^{\top}(QQ^{\top})^{-1}\mathbf{y}\leq\left(\dfrac{1}{L}\right)^{2}\right\}=\left\{\mathbf{y}:\left\lVert U\mathbf{y}\right\rVert\leq\left(\dfrac{1}{L}\right)\right\}$ where $U=(QQ^{\top})^{-1/2}$. By Lemma 5.36 and the definition of $\mathcal{L}_{s}(B^{(i)})$, we can find a $l_{2}$-ball with radius $\frac{\mathcal{L}_{s}(B^{(i)})}{L}$ that both contains $\mathbf{v}$ and is included in $E$ and therefore included in $\mathcal{P}^{(i)}$.
@@
The smoothness assumption in Theorem 5.38 can be relaxed. Instead of requiring that $f$ be globally smooth, that is, globally upper bounded by a convex quadratic function, it suffices to assume that $f$ is locally upper bounded by a convex quadratic function at each point $\mathbf{x}$. In particular, as long as this local smoothness condition guarantees the existence of an $l_{2}$ -ball with nontrivial radius, the conclusion of Theorem 5.38 remains valid.
@@ Acknowledgments
We would also like to thank Diego Cifuentes, Greg Blekherman, Santosh Vempala, and Kazuo Murota for the various wonderful discussions.